An isolated free electron is traveling through an electric field from some initial point where it's coulombic potential energy per unit charge (voltage) is 17KJ/C and velocity = 93Mm/s to some final point where it's coulombic potential energy per unit charge is 6KJ/C. Determine the charge in velocity of the electron . Neglect gravitational forces. What is the solution please.
1
Expert's answer
2019-04-30T10:40:02-0400

We have


"\phi_1=17000 (V); \phi_2 =6000 (V); v_1=93\cdot10^6 (m/s)"


And we need: light velocity; electron rest mass; electron charge

"c=3\cdot 0^8 (m/s); m_0=9.1\cdot 10^{-31} (kg); q=-1.6\cdot 10^{-19} (C)"

Comparing the speed of the electron with the speed of light

"\frac{93\cdot 10^6 }{3\cdot 10^8}\cdot 100 = 30"

We see 30 percents of the light velocity, Then this is a relativistic electron.

Kinetic energy in relativistic theory

"W=m_0c^2(\frac{1}{\sqrt{1-\frac{v^2}{c^2}}}-1)"


For initial condition kinetic energy of electron

"W_1=9.1\cdot 10^{-31}\cdot (3\cdot 10^8)^2\cdot(\frac{1}{\sqrt{1-\frac{(93\cdot 10^6)^2}{(3\cdot 10^8)^2}}}-1)=4.244*10^{-15} (J)"


When electron travelling trough electric field, the energy change for

"\Delta W=q(\phi_1-\phi_2)""\Delta W=-1.6\cdot 10^{-19} \cdot (17000-6000)=-1.76\cdot 10^{-15} (J)"

The kinetic energy decreases

"W_2=W_1-\Delta W""W_2=4.244\cdot 10^{-15} - 1.76 \cdot 10^{-15} = 2.484\cdot 10^{-15} (J)"

And electron velocity from kinetic energy equal

"v_2=c\sqrt{1-\frac{1}{(\frac{W_2}{m_0\cdot c^2}+1)^2}}"

"v_2=72.25\cdot 10^6 (m/s)"

Change velocity

"\Delta v=v_2-v_1""\Delta v= (72.25-93)\cdot 10^6=-20.75 (Mm/s)"

Answer: velocity decreases on 20.75 Mm/s